OpenStudy (anonymous):

if sin theta < 0 and tan theta < 0 then:

OpenStudy (anonymous):

you are in quadrant 4

OpenStudy (anonymous):

so would it be 270 degrees < theta < 360 degrees?

OpenStudy (anonymous):

because if tangent is negative and sine is negative, since \[\tan(x)=\frac{\sin(x)}{\cos(x)}\] that makes \(\cos(x)>0\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you person

OpenStudy (anonymous):

Hey, can you help me with some more math?